An 9 hour exposure to a sound intensity level of 95.0 dB may cause hearing damage. What energy in joules falls on a 0.750 cm diameter eardrum so exposed?
Answer #1

given
sound intebnsity level, beta = 95 dB
time, delta_t = 9 hours = 9*60*60 s

d = 0.75 cm = 0.0075 m
A = pi*d^2/4

= pi*0.0075^2/4

= 4.42*10^-5 m^2

let I is the intensity of the sound.

use, beta = 10*log(I/Io)

95 = 10*log(I/10^-12)

95 = log(I/10^-12)

10^9.5 = I/10^-12

I = 10^(9.5-12)

I = 3.16*10^-3 W/m^2

we know, I = Power/Area

Power = I*Area

energy incident = Power*time taken

= I*Area*delta_t

= 3.16*10^-3*4.42*10^-5*9*60*60

= 0.00452 J <<<<<<<<<<---------------------------Answer